# Targeted Immunisation Thresholds for the Contact Process on Power-Law Trees

### v1

@@ 6.2 Small truncation threshold
$M_{t}:=\left({3}/{2}\right)^{H_{t}}$ will be a supermartingale while $H_{t}$ is sufficiently large. We rewrite
$\begin{split}\mathcal{L}M_{t}&\leq\lambda|\partial S|\left(\left(\frac{3}{2}\right)^{H_{t}+1}-\left(\frac{3}{2}\right)^{H_{t}}\right)+\sum_{v\in\xi_{t}}\mathcal{L}^{(v)}M_{t}\\ &=\frac{\lambda|\partial S|}{2}\left(\frac{3}{2}\right)^{H_{t}}+\sum_{v\in\xi_{t}}\mathcal{L}^{(v)}M_{t}\end{split}$ (14)
where $\mathcal{L}^{(v)}$ is simply the generator of the Markov process of infections and recoveries by $v$. First we consider infected vertices $v$ at distance $\operatorname{d}(v,S)\geq 3$ and so with weight $w=(2\lambda)^{\operatorname{d}(v,S)-1}$,
$\begin{split}\mathcal{L}^{(v)}M_{t}&\leq\left(\frac{3}{2}\right)^{H_{t}-w}\left(\frac{\lambda}{8\lambda^{2}}\left(\left(\frac{3}{2}\right)^{w(1+2\lambda)}-\left(\frac{3}{2}\right)^{w}\right)+\lambda\left(\left(\frac{3}{2}\right)^{w(1+\frac{1}{2\lambda})}-\left(\frac{3}{2}\right)^{w}\right)+1-\left(\frac{3}{2}\right)^{w}\right)\\ &\sim\left(\frac{3}{2}\right)^{H_{t}}\left(\frac{1}{8\lambda}\cdot 2\lambda w\log\frac{3}{2}+\lambda\cdot\frac{w}{2\lambda}\log\frac{3}{2}-w\log\frac{3}{2}\right)=w\left(\frac{3}{2}\right)^{H_{t}}\left(-\frac{1}{4}\log\frac{3}{2}\right)\end{split}$
where we were able to apply $e^{x}\sim 1+x$ to simplify every term in the limit $\lambda\downarrow 0$. We have a different calculation when $\operatorname{d}(v,S)=2$, in which case
@@
$\begin{split}\mathcal{L}^{(v)}M_{t}\leq\sim\left(\frac{3}{2}\right)^{H_{t}}\left(\frac{1}{4}w\log\frac{3}{2}+\lambda\left(\left(\frac{3}{2}\right)^{\frac{w}{2\lambda}}-1\right)-w\log\frac{3}{2}\right)=w\left(\frac{3}{2}\right)^{H_{t}}\left(-\frac{3}{4}\log\frac{3}{2}+\frac{1}{4}\right),\end{split}$
and in the final case $\operatorname{d}(v,S)=1$ infections of parent in $S$ do not count for $H_{t}$ and so
@@
$\begin{split}\mathcal{L}^{(v)}M_{t}\leq\frac{1}{8\lambda}\left(\left(\frac{3}{2}\right)^{H_{t}+2\lambda}-\left(\frac{3}{2}\right)^{H_{t}}\right)+\left(\frac{3}{2}\right)^{H_{t}-1}-\left(\frac{3}{2}\right)^{H_{t}}\sim w\left(\frac{3}{2}\right)^{H_{t}}\left(\frac{1}{4}\log\frac{3}{2}-\frac{1}{3}\right).\end{split}$
@@
Recalling (14), we have
$\mathcal{L}\left(\frac{3}{2}\right)^{H_{t}}\leq\left(\frac{3}{2}\right)^{H_{t}}\left(\frac{\lambda|\partial S|}{2}-\frac{1}{20}H_{t}\right)$
negative as long as $H_{t}>10\lambda|\partial S|$. Defining
$\tau_{D}:=\inf\left\{t>\tau_{0}:H_{t}<10\lambda|\partial S|\right\},$
we find that $(M_{t\wedge\tau_{D}})_{t}$ is a supermartingale. This supermartingale is a pure jump process so we must be slightly careful with hitting of sets, but the still the jumps cannot be larger than a distance of $1$. We define stopping times
$\tau_{0}:=\inf\left\{t>0:H_{t}\geq 15\lambda|\partial S|\right\}$
$\tau_{U}:=\inf\left\{t>\tau_{0}:H_{t}>20\lambda|\partial S|\right\}$
and then apply optional stopping to deduce
$\gamma:=\mathbb{P}\left(\tau_{U}<\tau_{D}\right)\leq\frac{\left(\frac{3}{2}\right)^{15\lambda|\partial S|+1}-\left(\frac{3}{2}\right)^{10\lambda|\partial S|-1}}{\left(\frac{3}{2}\right)^{20\lambda|\partial S|}-\left(\frac{3}{2}\right)^{10\lambda|\partial S|}}\sim\left(\frac{3}{2}\right)^{-5\lambda|\partial S|+1}.$ (15)
It remains to lower bound the time which each downward exit takes. For this we use a sequence $(E_{i})_{i}$ of i.i.d. exponentials of mean $1$
@@
$Z_{t}+\tilde{Z}_{t}+\tilde{\tilde{Z}}_{t}=2\sum_{v\in\xi_{\tau_{0}}}w(v)\geq 30\lambda|\partial S|;$
by symmetry $Z_{t}$ has at least $1/3$ chance of being the joint largest of the three versions and hence $\mathbb{P}(Z_{t}\geq 10\lambda|\partial S|)\geq 1/3$. All in all, we have a number of downward exits before $H_{t}$ can exceed $20\lambda|\partial S|$ which is at least a Geometric of mean $-1+1/\gamma\gtrsim\left({3}/{2}\right)^{5\lambda|\partial S|-1}$. First note, by Markov’s inequality,
@@
$\begin{split}\mathbb{P}\left(\tau_{U}-\tau_{0}<7^{\lambda|\partial S|}\right)&\leq\mathbb{P}\left(\operatorname{Geom}\left(\gamma\right)\leq\frac{7^{1+\lambda|\partial S|}}{\log\frac{3}{2}}\right)+\mathbb{P}\left(\operatorname{Bin}\left(\left\lfloor\frac{7^{1+\lambda|\partial S|}}{\log\frac{3}{2}}\right\rfloor,\frac{1}{3}\right)\leq\frac{7^{\lambda|\partial S|}}{\log\frac{3}{2}}\right)\\ &\leq\gamma\frac{7^{\lambda|\partial S|}}{\frac{2}{3}\log\frac{3}{2}}+\gamma+\exp\left(-\frac{7^{1+\lambda|\partial S|}}{24\log\frac{3}{2}}\right)\\ &\leq\exp\left(-\frac{1}{100\lambda}\right)\end{split}$

### v2

@@ 6.2 Small truncation threshold
Given the larger distance which is here necessary to prevent local divergence of the CP-BRW, we just repeat the proof of Corollary 5.2 to see that the expected number of infections absorbed at $v$ has
$m(\mathcal{C},v)\leq 3e^{200\lambda^{2}|\partial\mathcal{C}|}(2\lambda)^{\lceil\delta\log(1/\lambda)\rceil-1}.$
Now recalling from (12) the bound $|\partial\mathcal{C}|\leq\frac{\delta\log^{2}\frac{1}{\lambda}}{300\lambda^{2}}$ (because we assumed for this corollary that the environment is good), we have:
$3e^{200\lambda^{2}|\partial\mathcal{C}|}(2\lambda)^{\lceil\delta\log(1/\lambda)\rceil-1}\leq 3e^{-\frac{1}{3}\delta\log^{2}\frac{1}{\lambda}+(1+\delta\log 2)\log\frac{1}{\lambda}}\leq e^{-\frac{1}{4}\delta\log^{2}\frac{1}{\lambda}}.$
By Corollary 6.6, on a good environment the probability that the infection ever reaches any given vertex $v$ of $\mathcal{C}^{+}$ at maximal distance $\lceil\delta\log(1/\lambda)\rceil$ from $\mathcal{C}$ is bounded by $e^{-\frac{1}{4}\delta\log^{2}\frac{1}{\lambda}}$. Recalling that this good environment has $|\mathcal{C}^{+}|\leq\frac{\delta}{400\rho}\mathbbm{t}^{1+\delta\log\mu}$, by the union bound the survival probability of the infection on this good environment is bounded by
$\left\lfloor\frac{\delta}{400\rho}\mathbbm{t}^{1+\delta\log\mu}\right\rfloor e^{-\frac{1}{4}\delta\log^{2}\frac{1}{\lambda}}$
Proof of Lemma 6.3.
First, by the monotonicity of the Galton-Watson tree it suffices to show the result on the tree with maximal root degree $\operatorname{d}(o)=\mathbbm{t}-1$. We then explore the ball around this root of radius $\left\lceil\delta\log\nicefrac{{1}}{{\lambda}}\right\rceil$ for other big vertices (with degree at least $\frac{1}{8\lambda^{2}}$), as well as the balls around paths between big vertices.
Continue iteratively to explore the ball of radius $\left\lceil\delta\log\nicefrac{{1}}{{\lambda}}\right\rceil$ around every big vertex until this construction terminates and we have explored the whole $\mathcal{C}$, revealing the big vertices and paths between them of length at most $\left\lceil\delta\log\nicefrac{{1}}{{\lambda}}\right\rceil$.
To obtain the desired bounds on $|\mathcal{B}|$ and $|\mathcal{C}^{+}|$, we combine two ingredients:
  • We first control the size of the ball of radius $\left\lceil\delta\log\nicefrac{{1}}{{\lambda}}\right\rceil$ around a given big vertex. Using [strayappendix, Theorem 1.1] with
    $n=\left\lceil\delta\log\frac{1}{\lambda}\right\rceil\qquad\text{and}\qquad x=\lambda^{-2-\frac{1}{2}\delta\log\mu},$
    we obtain that the probability that the boundary of a ball with root degree $\lfloor 1/8\lambda^{2}\rfloor$ contains more than
    $\mathbbm{t}^{1+\delta\log\mu+\frac{1}{2}\delta\log\mu}$
    vertices decays like a stretched exponential in $\lambda$. A big vertex and the path which discovered it produce together at most
    $\left\lceil\delta\log\frac{1}{\lambda}\right\rceil+\frac{\mathbbm{t}}{\lfloor 1/8\lambda^{2}\rfloor}\leq 9\rho\log^{2}\left(\nicefrac{{1}}{{\lambda}}\right)$
    such balls and hence by the union bound, a total contribution to $\mathcal{C}^{+}$ smaller than
    $9\rho\log^{2}\left(\nicefrac{{1}}{{\lambda}}\right)\mathbbm{t}^{1+\frac{3}{4}\delta\log\mu}\leq\mathbbm{t}^{1+\frac{4}{5}\delta\log\mu}$
    with a probability that’s also stretched exponential in $\lambda$, and so in particular a probability that is smaller than any polynomial in $\lambda$.
    By the union bound again, we get that with a probability which is still smaller than any polynomial in $\lambda$, the union of the balls of radius $\left\lceil\delta\log\frac{1}{\lambda}\right\rceil$ around the first $\left\lfloor\frac{\delta}{400\rho}\right\rfloor$ explored big vertices and paths between them contain less than $\left\lfloor\frac{\delta}{400\rho}\mathbbm{t}^{1+\delta\log\mu}\right\rfloor$ vertices.
  • It remains to bound $|\mathcal{B}|$ by $\frac{\delta}{400\rho}$. To this end, it suffices to check that above $\left\lfloor\frac{\delta}{400\rho}\mathbbm{t}^{1+\delta\log\mu}\right\rfloor$ explored vertices, it is likely to find less than $\frac{\delta}{400\rho}$ big vertices. Note that for small enough $\lambda$ each explored vertex is big with probability
    $\mathbb{P}\left(D\geq\frac{1}{8\lambda^{2}}\right)<\mathbb{P}\left(X\geq\frac{1}{8\lambda^{2}}\right)<c_{X}\lambda^{2\tau-4}$
    and hence in $\left\lfloor\frac{\delta}{400\rho}\mathbbm{t}^{1+\delta\log\mu}\right\rfloor$ explored vertices we find fewer than $\frac{\delta}{400\rho}$ large vertices with probability, using the usual Chernoff bound, bounded by
    $\mathbb{P}\left(\operatorname{Bin}\left(\left\lfloor\frac{\delta}{400\rho}\mathbbm{t}^{1+\delta\log\mu}\right\rfloor,c_{X}\lambda^{2\tau-4}\right)\geq\frac{\delta}{400\rho}\right)$
@@
    $\begin{split}&\leq\mathbb{P}\left(\operatorname{Pois}\left(\frac{\delta}{400\rho}\rho^{1+\delta\log\mu}c_{X}\lambda^{2\tau-6-2\delta\log\mu}\log^{2+2\delta\log\mu}\frac{1}{\lambda}\right)\geq\frac{\delta}{400\rho}\right)\\ &\leq\lambda^{\frac{\varphi\delta}{400\rho}}\exp\left(\frac{\delta}{400\rho}\rho^{1+\delta\log\mu}c_{X}\lambda^{2\tau-6-2\delta\log\mu-\varphi}\log^{2+2\delta\log\mu}\frac{1}{\lambda}\right)\sim\lambda^{\frac{\varphi\delta}{400\rho}}\leq\lambda^{C}\end{split}$
    where we have set parameter $\varphi={400\rho C}/{\delta}$ so that by taking $\delta$ small and then $\rho$ small enough we guarantee $02\tau-6-2\delta\log\mu-\varphi>$ 0.
Proof of Lemma 6.4.
$M_{t}:=\left({3}/{2}\right)^{H_{t}}$ will be a supermartingale while $H_{t}$ is sufficiently large. Indeed, write its expected infinitesimal increase as
@@
$\begin{split}\frac{\operatorname{\mathbb{E}}\left[M_{t+\rm dt}-M_{t}\Big{|}\mathcal{F}_{t}\right]}{\rm dt}&=\lambda|\partial\mathcal{C}|\left(\left(\frac{3}{2}\right)^{H_{t}+1}-\left(\frac{3}{2}\right)^{H_{t}}\right)+\sum_{v\in\xi_{t}}\mathcal{L}(v)\\ &=\frac{\lambda|\partial\mathcal{C}|}{2}\left(\frac{3}{2}\right)^{H_{t}}+\sum_{v\in\xi_{t}}\mathcal{L}(v),\end{split}$ (15)
where again $\mathcal{F}_{t}$ is the $\sigma$-algebra generated by $(H_{s})_{s\leq t}$, and $\mathcal{L}(v)$ is the contribution to the expected infinitesimal increase due to infections and recoveries by $v$.
First we consider infected vertices $v$ at distance $\operatorname{d}(v,\mathcal{C})\geq 3$ and so with weight $w=(2\lambda)^{\operatorname{d}(v,\mathcal{C})-1}$,
$\begin{split}\mathcal{L}(v)&\leq\left(\frac{3}{2}\right)^{H_{t}-w}\left(\frac{\lambda}{8\lambda^{2}}\left(\left(\frac{3}{2}\right)^{w(1+2\lambda)}-\left(\frac{3}{2}\right)^{w}\right)+\lambda\left(\left(\frac{3}{2}\right)^{w(1+\frac{1}{2\lambda})}-\left(\frac{3}{2}\right)^{w}\right)+1-\left(\frac{3}{2}\right)^{w}\right)\\ &\sim\left(\frac{3}{2}\right)^{H_{t}}\left(\frac{1}{8\lambda}\cdot 2\lambda w\log\frac{3}{2}+\lambda\cdot\frac{w}{2\lambda}\log\frac{3}{2}-w\log\frac{3}{2}\right)=w\left(\frac{3}{2}\right)^{H_{t}}\left(-\frac{1}{4}\log\frac{3}{2}\right)\end{split}$
where we were able to apply $e^{x}\sim 1+x$ to simplify every term in the limit $\lambda\downarrow 0$. We have a different calculation when $\operatorname{d}(v,\mathcal{C})=2$, in which case
$\begin{split}\mathcal{L}(v)&\leq(1+o(1))\left(\frac{3}{2}\right)^{H_{t}}\left(\frac{1}{4}w\log\frac{3}{2}+\lambda\left(\left(\frac{3}{2}\right)^{\frac{w}{2\lambda}}-1\right)-w\log\frac{3}{2}\right)\\ &=w\left(\frac{3}{2}\right)^{H_{t}}\left(-\frac{3}{4}\log\frac{3}{2}+\frac{1}{4}\right),\end{split}$
and in the final case $\operatorname{d}(v,\mathcal{C})=1$ infections of parent in $\mathcal{C}$ do not count for $H_{t}$ and so
$\begin{split}\mathcal{L}(v)\leq\frac{1}{8\lambda}\left(\left(\frac{3}{2}\right)^{H_{t}+2\lambda}-\left(\frac{3}{2}\right)^{H_{t}}\right)+\left(\frac{3}{2}\right)^{H_{t}-1}-\left(\frac{3}{2}\right)^{H_{t}}\sim w\left(\frac{3}{2}\right)^{H_{t}}\left(\frac{1}{4}\log\frac{3}{2}-\frac{1}{3}\right).\end{split}$
@@
Recalling (15), we have that the expected infinitesimal increase of $M$ is bounded by
$\left(\frac{3}{2}\right)^{H_{t}}\left(\frac{\lambda|\partial\mathcal{C}|}{2}-\frac{1}{20}H_{t}\right)$
and thus negative as long as $H_{t}>10\lambda|\partial\mathcal{C}|$. Define
$\tau_{0}$ $:=\inf\left\{t\geq 0:H_{t}\geq 15\lambda|\partial\mathcal{C}|\right\},$
$\tau_{D}$ $:=\inf\left\{t>\tau_{0}:H_{t}<10\lambda|\partial\mathcal{C}|\right\},$
$\tau_{U}$ $:=\inf\left\{t>\tau_{0}:H_{t}>20\lambda|\partial\mathcal{C}|\right\}.$
On $[\tau_{0},\tau_{D}]$, the expected infinitesimal increase of $M$ is negative, and we have $H_{\tau_{0}}\leq 15\lambda|\partial\mathcal{C}|+1$ as well as $H_{\tau_{D}}\geq 10\lambda|\partial\mathcal{C}|-1$, so we can apply optional stopping to deduce
$\gamma:=\mathbb{P}\left(\tau_{U}<\tau_{D}\right)\leq\frac{\left(\nicefrac{{3}}{{2}}\right)^{15\lambda|\partial\mathcal{C}|+1}-\left(\nicefrac{{3}}{{2}}\right)^{10\lambda|\partial\mathcal{C}|-1}}{\left(\nicefrac{{3}}{{2}}\right)^{20\lambda|\partial\mathcal{C}|}-\left(\nicefrac{{3}}{{2}}\right)^{10\lambda|\partial\mathcal{C}|}}\sim\left(\frac{3}{2}\right)^{-5\lambda|\partial\mathcal{C}|+1}.$ (16)
We will thus typically observe many downward exits from above $15\lambda|\partial\mathcal{C}|$ to below $10\lambda|\partial\mathcal{C}|$ before $H_{t}$ can exceed $20\lambda|\partial\mathcal{C}|$, and it remains to lower bound the time which each downward exit takes. From a particular infection state, the contact process dominates the contact process with no infection allowed and only recovery. Hence, introducing a sequence $(E_{i})_{i}$ of i.i.d. exponentials of mean $1$, we have the following stochastic lower bound
@@
$Z_{t}+\tilde{Z}_{t}+\tilde{\tilde{Z}}_{t}=2\sum_{v\in\xi_{\tau_{0}}}w(v)\geq 30\lambda|\partial\mathcal{C}|.$
In particular, the sum of probabilities $\mathbb{P}(Z_{t}\geq 10\lambda|\partial\mathcal{C}|)+\mathbb{P}(\tilde{Z}_{t}\geq 10\lambda|\partial\mathcal{C}|)+\mathbb{P}(\tilde{\tilde{Z}}_{t}\geq 10\lambda|\partial\mathcal{C}|)$ exceeds one, but as these probabilities are equal, we deduce $\mathbb{P}(Z_{t}\geq 10\lambda|\partial\mathcal{C}|)\geq 1/3$.
All in all, we have a number of downward exits before $H_{t}$ can exceed $20\lambda|\partial\mathcal{C}|$ which is at least a Geometric of mean $-1+1/\gamma\gtrsim\left({3}/{2}\right)^{5\lambda|\partial\mathcal{C}|-1}$. First note, by Markov’s inequality,
@@
$\begin{split}\mathbb{P}\left(\tau_{U}-\tau_{0}<7^{\lambda|\partial\mathcal{C}|}\right)&\leq\mathbb{P}\left(\operatorname{Geom}\left(\gamma\right)\leq\frac{7^{1+\lambda|\partial\mathcal{C}|}}{\log\frac{3}{2}}\right)+\mathbb{P}\left(\operatorname{Bin}\left(\left\lfloor\frac{7^{1+\lambda|\partial\mathcal{C}|}}{\log\frac{3}{2}}\right\rfloor,\frac{1}{3}\right)\leq\frac{7^{\lambda|\partial\mathcal{C}|}}{\log\frac{3}{2}}\right)\\ &\leq\gamma\frac{7^{\lambda|\partial\mathcal{C}|}}{\frac{2}{3}\log\frac{3}{2}}+\gamma+\exp\left(-\frac{7^{1+\lambda|\partial\mathcal{C}|}}{24\log\frac{3}{2}}\right)\\ &\leq\exp\left(-\frac{1}{100\lambda}\right)\end{split}$